OpenStudy (anonymous):

what is the exact volume of the largest cone that can fit into a cube with edges of 10 inches?

OpenStudy (jack1):

cube volume = L x L x L = L^3 cone volume = pi x r^2 x h/3 in this case: r = L/2 and h = L cone volume = pi x (L/2)^2 x L/3 cone volume = pi x L^2/4 x L/3 cone volume = pi x L^3/12 so if L = 10 cube volume = L x L x L = L^3 cube volume = 10 x 10 x 10 = 10^3 cube volume = 1000 in^3 so if L = 10 cone volume = pi x r^2 x h/3 cone volume = pi x L^3/12 cone volume = pi x 10^3/12 cone volume = pi x 1000/12 cone volume = 261.8 ...ish in^3

OpenStudy (anonymous):

Thanks, man for your help.

OpenStudy (jack1):

no worries dude